OpenStudy (anonymous):

what two perfect squares gives you a sum of 58

OpenStudy (kendricklamar2014):

58 / 2 = ?

OpenStudy (kendricklamar2014):

@nyclovee

OpenStudy (anonymous):

58/2 isnt a whole number

OpenStudy (anonymous):

nvm it is

OpenStudy (kendricklamar2014):

divide 58 / 2 and that will give u ur answer

OpenStudy (anonymous):

29

OpenStudy (kendricklamar2014):

Yes ^ that is ur answer

OpenStudy (anonymous):

No you need two perfect sqaures

OpenStudy (anonymous):

no it's 49 and 9, they are both perfect squares

OpenStudy (anonymous):

meaning a number can multiply by itself to get to that number, example 3*3=9, 4*4=16, 25*25=625. 625, 16, and 9 are all perfect squares.